# Spatial vertical regression for spatial panel data: Evaluating the effect of the Florentine tramway’s first line on commercial vitality

### v2

@@ 4.1 The setup
We measure an outcome on the geography of interest for each time period, denoted as $\mathcal{Y}_{t}=\{\mathcal{Y}_{t}(\omega),\ \omega\in\Omega\}$, and $t=1,2,\dots,T$. The outcome is compacted over the areas $A_{i}$ to represent an overall outcome in the area at time $t$, denoted as $Y_{it}$. For example, if $\mathcal{Y}_{t}$ is an integrable surface on $\Omega$, $Y_{it}$ might represent the total value over the set $A_{i}$, defined as $Y_{it}=\int_{A_{i}}\mathcal{Y}_{t}(\omega)\mathrm{d}\omega$. In the case where $\mathcal{Y}_{t}$ is binary over $\Omega$, the outcome of area $Y_{it}$ might represent the number of locations in $A_{i}$ for which $\mathcal{Y}_{t}$ is equal to 1 at time $t$.
In our study, $\Omega$ corresponds to a large geographical area corresponding to the municipality of Florence, which we follow during the years 1996 to 2014. The set $S$ denotes the set of coordinates of the Florentine tramway stops, for which construction started in 2006. Therefore, we follow the area of Florence before the initiation of construction, and after construction has begun. The $N_{1}$ units correspond to areas surrounding the tramway stops that are at most at $d_{i}$ walking time distance away from a tramway stop, whereas the remaining $N_{0}$ units correspond to areas that are far away from the locations of tramway construction, as shown in Figure 1. Then, the spatial outcome $\mathcal{Y}_{t}$ is a binary indicator for each location denoting the presence or absence of a business at location $\omega$ at year $t$, and $Y_{it}$ denotes the number of businesses in area $A_{i}$ during the same year.
@@ 4.2 Causal estimands and assumptions
1 formalizes that none of the areas responded to the treatment before the treatment was actually initiated, and that the observed outcomes before the intervention time $T_{0}+1$ are not affected by whether treatment will be initiated or not. This assumption is often referred to as the non-anticipation of the treatment assumption (e.g. Bojinov and Shephard, 2019; Menchetti and Bojinov, 2020). Therefore, it is important that $T_{0}$ is decided based on the earlier time of the line’s construction phase, rather than based on its time of operation (Papadogeorgou et al., 2023). 2 states that the $N_{0}$ areas that are far from the locations in $S$ are always unaffected by whether these locations receive the treatment or not. Therefore, under 2, the $N_{0}$ areas $A_{i}$, for $i=N_{1}+1,N_{1}+2,\dots,N$ are referred to as the control areas/units and the $N_{1}$ areas $A_{i}$, for $i=1,2,\dots,N_{1}$, are referred to as the treated areas/units. Since the control areas are unaffected by the treatment initiation, they form a valid donor pool for imputing the missing control potential outcomes for the treated areas.
@@ 5 Bayesian approach for Spatial Vertical Regression
5.1 Formalization within the Bayesian paradigm
@@
The parameters of the linear combinations are the intercepts $\beta_{0}^{(1)},\ldots,\beta_{0}^{(N_{1})}$ and the weights for each control unit across the regressions of all treated units, $\bm{\beta}_{c}\equiv\left(\beta_{c}^{(1)},\ldots,\beta_{c}^{(N_{1})}\right)$, $c=N_{1}+1,\ldots,N_{1}+N_{0}$. These coefficients are chosen using a Bayesian regression approach, where the pre-treatment outcomes of the treated units are regressed on the pre-treatment outcomes of the control units. Formally, for $i=1,2,\ldots,N_{1}$,
@@
In Equation Equation 6 the observed data in the pre-intervention period are used to estimate the model parameters for the potential outcomes model in Equation 5. We specify that the errors $\epsilon_{it}$ are zero mean Normal error terms possibly spatially correlated: $\bm{\epsilon}_{t}=(\epsilon_{1t},\ldots,\epsilon_{N_{1}t})^{\top}\sim\text{MVN}_{N_{1}}(\bm{0}_{N_{1}},\Sigma_{e}),$ where $\text{MVN}_{N_{1}}(\bm{0}_{N_{1}},\Sigma_{e})$ denotes a $N_{1}-$multivariate Normal distribution with vector mean zero and covariance matrix $\Sigma_{e}$. We discuss the choice of $\Sigma_{e}$ below.
@@ 5.2.1 Gaussian process specification on control unit’s regression coefficients
$\bm{\beta}_{c}\sim\mathcal{GP}(b_{c}\mathbf{1}_{N_{1}},\Sigma_{\beta})$ (7)
with $b_{c}\sim\mathcal{N}(\mu_{b},\sigma^{2}_{b})$ and $\mathbf{1}_{N_{1}}$ the $N_{1}-$dimensional vector with all elements equal to one. Importantly, the coefficients’ covariance matrix is set to
@@ 5.2.2 Spatial dependence in outcome model residuals
The Bayesian regression approach we propose presents several advantages. We embed information on the spatial correlation structure across units directly into the model. Through the Gaussian Process priors on the weights of the linear combinations defining the control potential outcomes for the treated units, we inform the model that those weights are expected to vary smoothly across treated units whose distance from the treatment sites increases progressively. This reflects our belief that the control potential outcomes for treated units that are proximal in space are expected to be similar, and the similarity diminishes for treated units that are located far away. Although the information on the spatial correlation structure enters the regression coefficients, we also allow for residual spatial correlation. Last but not least, the Bayesian regression approach naturally provides a quantification of the uncertainty of the imputed control potential outcomes for the treated units, and thus, of the causal effects. This is an important advantage. The quantification of the uncertainty in causal panel studies, especially when SC methods are used, is still an open issue and subject to debate (e.g., Shen et al., 2022).
The importance of accounting for dependence across units, time and outcomes in causal panel-data studies has recently attracted the attention of the causal inference community (e.g., Ben-Michael et al., 2023; Sun et al., 2023). Our method presents some similarities with the method recently proposed by Ben-Michael et al. (2023): we both use Gaussian processes to account for dependence across units, but in a different way. Ben-Michael et al. (2023) assume that the model for the control outcomes for all units is a unit-specific linear combination of latent time-varying variables shared across units. They specify a Gaussian process for each vector of time-varying latent variables, where the kernel of the Gaussian process captures the correlation of the latent variables across time. Instead, we embed the spatial structure information in the Gaussian Process priors of the coefficients of the control units in the linear combinations which define the control potential outcomes for the treated units. The kernels of the Gaussian Process priors impose a structure across the unit-specific weights of the treated units that reflects the spatial correlation structure across them.
@@ 5.3 Choice of prior distributions and computation
We impose the following priors on the hyperparameters $\sigma^{2}_{\beta}$ and $\rho^{2}_{\beta}$:
$\sigma^{2}_{\beta}\sim\hbox{Inverse-Gamma}(a_{\beta},\eta_{\beta})\qquad\text{and}\qquad\rho^{2}_{\beta}\sim\hbox{Gamma}(a_{\rho_{\beta}},\eta_{\rho_{\beta}}),$
where the Gamma distribution is specified using the shape and rate parameters, and the Inverse-Gamma is specified using the shape and scale parameters. For the remaining parameters of the model, the intercepts, $\beta_{0}^{(i)}$, $i=1,\ldots,N_{1}$; the overall variance of the error terms, $\sigma^{2}_{e}$; and the parameters $\rho_{e}$ and $w$ governing the errors’ correlation structure, we specify the following priors:
$\beta_{0}^{(i)}\sim N(\mu_{0},\sigma^{2}_{0})\quad i=1,\ldots,N_{1},$
$\sigma^{2}_{e}\sim\hbox{Inverse-Gamma}(a_{e},\eta_{e}),$
@@
$w\sim\hbox{Uniform}(0,1).$
Disparities in outcome magnitudes across units can affect the estimation of spatially correlated coefficients, leading the model to focus on absolute levels rather than on the underlying trends. We bypass this challenge by normalizing the data which corrects for differences in the overall level and size across units. We normalize the time series of each treated and control units’ outcomes with respect to the pre-treatment mean and standard deviation (as suggested also in Abadie et al., 2010; Abadie, 2021). Specifically, we use
@@
where $\overline{Y}_{i}=\sum_{t=1}^{T_{0}}Y_{it}/T_{0}$. In what follows, we assume that the outcomes $Y$ are normalized, and avoid using the notation $\widetilde{Y}$.
Working with normalized data allows us to specify values for hyperparameters that are expected to perform well across a variety of settings (as also evident in our simulations in Section 6). Along with standardizing the outcomes, we rescale the values $d_{i}$ to vary from 0 to 1. Since employing diffuse, non-informative prior distributions can perform poorly (Gelman et al., 2008), we instead adopt weakly informative prior distributions. For normalized data we set the hyperparameters of the prior distributions as follows: $\mu_{b}=0$, $\sigma^{2}_{b}=1$, $\mu_{0}=0$, $\sigma^{2}_{0}=1$, $a_{\beta}=2.7$, $\eta_{\beta}=0.3$, $a_{\rho_{\beta}}=0.5$, $\eta_{\rho_{\beta}}=2$, $a_{e}=2$, $\eta_{e}=0.35$, $a_{\rho_{e}}=0.5$, and $\eta_{\rho_{e}}=1.5$. These choices are adopted throughout our work. A formal justification of these choices along with draws from the induced prior distributions for $\bm{\beta}_{c}$ and the residual spatial correlation of treated areas are given in Supplement 1.
The posterior distribution of the model parameters and the missing control potential outcomes for the treated units is derived using Hamiltonian Monte Carlo (HMC) implemented in Stan (Carpenter et al., 2017) and its R interface (Stan Development Team, 2018). The code developed to obtain the results presented in this paper is publicly available at the following GitHub repository: https://github.com/Giulio-Grossi/SVR. This software enables the replication of all experiments described in the paper in a spatial context and, more generally, can be employed in any setting where a generic distance measure between units is defined. As a result, the proposed method is highly flexible and can be adapted to a variety of applications and analytical contexts. Furthermore, the repository includes the code for obtaining estimates and confidence intervals using the other methods evaluated in the simulations of Section 6.
@@ 6.1 Data generating process
In this section, we present the simulation study we conducted to investigate the extent to which the spatial correlation across the treated units can influence the regression coefficients and to evaluate the performance of our approach with respect to the original SC method and a standard vertical regression approach under different scenarios, including a scenario without spatial structure.
We consider a causal panel study with $N_{1}=5$ treated areas and $N_{0}=10$ control areas observed in $T_{0}$ pre-treatment periods and $T-T_{0}$ post-treatment periods. We consider three lengths of the pre-treatment period, $T_{0}=10,20,40$, combined with the following lengths of the post-treatment period: $T-T_{0}=5,10,T_{0}/2$, for a total of seven distinct scenarios (for $T_{0}=10$, $T-T_{0}=T_{0}/2=5$, and for $T_{0}=20$, $T-T_{0}=T_{0}/2=10$). Therefore, we consider scenarios that resemble our data described in Section 3.
@@
We set $\sigma^{2}_{s}=0.4$ and consider four values of the lengthscale parameter $\rho_{s}^{2}$: $0.001^{2}$, $0.2^{2}$, $0.4^{2}$, and $0.6^{2}$ corresponding to scenarios where the weights associated with the same control unit for different treated units are essentially independent, weakly correlated, mid-correlated, and strongly correlated, respectively. These scenarios mimic studies where accounting for the spatial correlation structure among the treated units in estimating the regression weights may be progressively more relevant, and the case of $\rho_{s}^{2}=0.001^{2}$ represents a scenario where our approach is not needed. In Supplement 2, we include a visualization of $\bm{\beta}_{c}$ in our simulations under the different values of $\rho_{s}^{2}$.
We generate outcomes for control areas that are temporally correlated drawing from temporal Gaussian processes. Specifically, let $\bm{Y}_{c}^{\bm{z}_{0}}=\bm{Y}_{c}=(Y_{c1},\ldots,Y_{cT})^{\top}$ represent the times-series under no intervention for the $c^{th}$ unit for $c=N_{1}+1,\ldots,N_{1}+N_{0}$, which we observe. We draw $\bm{Y}_{c}$ from a $T-$ dimensional multivariate normal distribution. We specify mean vector $t_{c}\bm{1}_{T}$ where $t_{c}\sim\mathcal{N}(0,0.7^{2})$. The covariance matrix is set to $\Sigma_{T\times T}=\sigma^{2}_{c}\mathcal{K}(D_{T},\rho^{2}_{c})+\eta_{c}\mathbf{I}_{T}$, where $D_{T}$ is a $T\times T$ matrix of time differences and $\mathbf{I}_{T}$ is the $T\times T$ identity matrix, so that
$\operatorname{Cov}(Y_{ct},Y_{ct^{\prime}})=\sigma^{2}_{c}\operatorname{exp}\left(-\frac{[D_{T}]_{tt^{\prime}}}{2\rho_{c}^{2}}\right)+\eta_{c}\text{I}(t=t^{\prime}),$
where $\mathrm{I}\{\cdot\}$ is the indicator function. We set $\sigma^{2}_{c}=0.3^{2}$, $\rho^{2}_{c}=0.05^{2}$, $\eta_{c}=0.15^{2}$, for all $c=N_{1}+1,\ldots,N_{1}+N_{0}$. In Supplement 2, we illustrate that the outcome time series for control units under this design has temporal correlation that resembles the data in Figure 2.

### v3

@@ 4.1 The setup
We measure an outcome on the geography of interest for each time period, denoted as $\mathcal{Y}_{t}=\{\mathcal{Y}_{t}(\omega),\ \omega\in\Omega\}$, and $t=1,2,\dots,T$. The outcome is compacted over the areas $A_{i}$ to represent an overall outcome in the area at time $t$, denoted as $Y_{it}$. For example, if $\mathcal{Y}_{t}$ is an integrable surface on $\Omega$, $Y_{it}$ might represent the total value over the set $A_{i}$, defined as $Y_{it}=\int_{A_{i}}\mathcal{Y}_{t}(\omega)\mathrm{d}\omega$. In the case where $\mathcal{Y}_{t}$ is binary over $\Omega$, the outcome $Y_{it}$ might instead represent the number of locations in $A_{i}$ for which $\mathcal{Y}_{t}$ is equal to 1 at time $t$.
In our study, $\Omega$ corresponds to the geographical area of the municipality of Florence, which we follow during the years 1996 to 2016. The set $S$ denotes the set of coordinates of the Florentine tramway stops, for which construction started in 2006. Therefore, we follow the area of Florence before the initiation of construction, and after construction has begun. The $N_{1}$ units correspond to areas surrounding the tramway stops that are within $d_{i}$ walking time distance from a tramway stop, whereas the remaining $N_{0}$ units correspond to areas that are far away from the locations of the tramway, as shown in Figure 1. Then, the spatial outcome $\mathcal{Y}_{t}$ is a binary indicator for each location denoting the presence or absence of a business at location $\omega$ at year $t$, and $Y_{it}$ denotes the number of businesses in area $A_{i}$ during the same year.
@@ 4.2 Causal estimands and assumptions
1 formalizes that none of the areas responded to the treatment before the treatment was actually initiated, and that the observed outcomes before the intervention time $T_{0}+1$ are not affected by whether treatment will be initiated or not. This assumption is often referred to as the non-anticipation of the treatment assumption (e.g. Bojinov and Shephard, 2019; Menchetti and Bojinov, 2020).
For that reason, we set $T_{0}$ based on the initiation of the line’s construction phase, rather than based on its time of operation (Papadogeorgou et al., 2023). We believe that this choice renders 1 plausible in our setting, as it is unlikely that the tramway line affected outcomes prior to its construction phase. The reason is that, even though the light rail project in Florence was announced in 2000, its implementation was marked by repeated failed tenders, prolonged legal disputes, and substantial public uncertainty regarding its feasibility and timing. Given this highly unstable and contested process, it is unlikely that private economic agents, in particular store owners, engaged in systematic anticipatory behaviour prior to the actual construction phase.
2 states that the $N_{0}$ areas that are far from the locations in $S$ are always unaffected by whether these locations receive the treatment or not. Therefore, under 2, the $N_{0}$ areas $A_{i}$, for $i=N_{1}+1,N_{1}+2,\dots,N$ are referred to as the control areas/units and the $N_{1}$ areas $A_{i}$, for $i=1,2,\dots,N_{1}$, are referred to as the treated areas/units. Since the control areas are unaffected by the treatment initiation, they form a valid donor pool for imputing the missing control potential outcomes for the treated areas.
@@ 5 Bayesian approach for Spatial Vertical Regression
5.1 Formalization of vertical regression within the Bayesian paradigm
@@
The parameters of the linear combinations are the intercepts $\beta_{0}^{(1)},\ldots,\beta_{0}^{(N_{1})}$ and the weights for each control unit across the regressions of all treated units, $\bm{\beta}_{c}\equiv\left(\beta_{c}^{(1)},\ldots,\beta_{c}^{(N_{1})}\right)$, $c=N_{1}+1,\ldots,N_{1}+N_{0}$. The model in Equation 5 on the potential outcomes is the crucial assumption for estimation, as also employed in Abadie et al. (2010) and Athey et al. (2021). Therefore, we assume stable pattern between pre and post-treatment outcomes which implies that the conditional mean of the potential outcomes can be written in the form of Equation 5 with fixed coefficients across time. We also assume the absence of idiosyncratic shocks in that the distribution of the error terms is unaltered across time, as in Abadie et al. (2015).
These coefficients are chosen using a Bayesian regression approach, where the pre-treatment outcomes of the treated units are regressed on the pre-treatment outcomes of the control units. Formally, for $i=1,2,\ldots,N_{1}$,
@@
In Equation Equation 6 the observed data in the pre-intervention period are used to estimate the model parameters for the potential outcomes model in Equation 5. We specify that the $\epsilon_{it}$ are zero mean Normal error terms, possibly spatially correlated, $\bm{\epsilon}_{t}=(\epsilon_{1t},\ldots,\epsilon_{N_{1}t})^{\top}\sim\text{MVN}_{N_{1}}(\bm{0}_{N_{1}},\Sigma_{e}),$ where $\text{MVN}_{N_{1}}(\bm{0}_{N_{1}},\Sigma_{e})$ denotes a $N_{1}$-multivariate Normal distribution with vector mean zero and covariance matrix $\Sigma_{e}$. We discuss the choice of $\Sigma_{e}$ below.
@@ 5.2.1 Gaussian process specification on control unit’s regression coefficients
$\bm{\beta}_{c}\sim\mathcal{GP}(b_{c}\mathbf{1}_{N_{1}},\Sigma_{\beta}),$ (7)
where $b_{c}\mathbf{1}_{N_{1}}$ is the vector of means for the Gaussian process with $b_{c}\in\mathbb{R}$ a scalar parameter and $\mathbf{1}_{N_{1}}$ the $N_{1}$-dimensional vector with all elements equal to one. The coefficients’ covariance matrix is set to
@@ 5.2.2 Spatial dependence in outcome model residuals
The Bayesian regression approach we propose presents several advantages. We embed information on the spatial correlation structure across units directly into the model. Through the Gaussian Process priors on the weights of the linear combinations defining the control potential outcomes for the treated units, we inform the model that those weights are expected to vary smoothly across treated units whose distance from the treatment sites increases progressively. This reflects our belief that nearby treated units have similar relationships to control units, and the similarity diminishes as the distance of the treated units increases. Additionally, we allow for spatial correlation in the outcome model residuals. Last but not least, the Bayesian regression approach naturally provides a quantification of the uncertainty of the imputed control potential outcomes for the treated units, and thus, of the causal effects. This is an important advantage. The quantification of the uncertainty in causal panel studies, especially when SC methods are used, is still an open issue and subject to debate (Shen et al., 2022).
The importance of accounting for dependence across units, time and outcomes in causal panel-data studies has recently attracted the attention of the causal inference community (e.g., Ben-Michael et al., 2023; Sun et al., 2023). Our method presents some similarities with the method recently proposed by Ben-Michael et al. (2023). We both use Gaussian processes to account for dependence across units, but in a different way. Ben-Michael et al. (2023) assume that the model for the control outcomes for all units is a unit-specific linear combination of latent time-varying variables shared across units. They specify a Gaussian process for each vector of time-varying latent variables, where the kernel of the Gaussian process captures the correlation of the latent variables across time. Instead, we embed the spatial structure information in the Gaussian Process priors of the coefficients of the control units in the linear combinations which define the control potential outcomes for the treated units. The kernels of the Gaussian Process priors impose a structure across the unit-specific weights of the treated units that reflects the spatial correlation structure across them.
@@ 5.3 Choice of prior distributions and computation
First, we discuss priors that relate to the control units’ coefficients. To encourage regularization in low-information settings such as data with a small number of pre-intervention time periods, we specify a Laplace prior on a control unit’s overall importance across treated areas, $b_{c}\sim\hbox{Laplace}(\lambda_{b})$. We adopt prior distributions $\sigma^{2}_{\beta}\sim\mathcal{N^{+}}(a_{\beta},\eta_{\beta})$ and $\rho^{2}_{\beta}\sim\hbox{Gamma}(a_{\rho_{\beta}},\eta_{\rho_{\beta}}),$ where the Gamma distribution is specified using the shape and rate parameters, and $\mathcal{N^{+}}$ represents the Normal distribution truncated below at zero.
For the remaining parameters of the model, including the intercepts, the variance of the residual error terms, and the parameters governing the errors’ correlation structure, we specify the following priors:
$\beta_{0}^{(i)}\sim N(\mu_{0},\sigma^{2}_{0}),\quad i=1,\ldots,N_{1},$
$\sigma^{2}_{e}\sim\mathcal{N^{+}}(a_{e},\eta_{e}),$
@@
$w\sim\hbox{Beta}(a_{w},b_{w}).$
Disparities in outcome magnitudes across units can affect the estimation of spatially correlated coefficients, leading the model to focus on absolute levels rather than on the underlying trends. We bypass this challenge by normalizing the data to correct for differences in the overall level and size across units. We normalize the time series of each treated and control units’ outcomes with respect to the pre-treatment mean and standard deviation (as suggested also in Abadie et al., 2010; Abadie, 2021). Specifically, we use
@@
where $\overline{Y}_{i}=\sum_{t=1}^{T_{0}}Y_{it}/T_{0}$. In what follows, we assume that the outcomes $Y$ are normalized, and avoid using the notation $\widetilde{Y}$. Along with standardizing the outcomes, we rescale the values $d_{i}$ to vary from 0 to 1.
Working with normalized data allows us to specify values for hyperparameters that are expected to perform well across a variety of settings (as also evident in our simulations in Section 6). Since employing diffuse, non-informative prior distributions can perform poorly (Gelman et al., 2008), we instead adopt weakly informative prior distributions. For normalized data, we set the hyperparameters of the prior distributions as follows: $\lambda_{b}=0.1$ $\mu_{0}=0$, $\sigma^{2}_{0}=1$, $a_{\beta}=0$, $\eta_{\beta}=0.35$, $a_{\rho_{\beta}}=0.5$, $\eta_{\rho_{\beta}}=2$, $a_{e}=0$, $\eta_{e}=0.5$, $a_{\rho_{e}}=0.5$, $\eta_{\rho_{e}}=1.5$, and $a_{w}=2,\text{ }b_{w}=2$. These choices are adopted throughout our work. A formal justification of these choices along with draws from the induced prior distributions for $\bm{\beta}_{c}$ and the residual spatial correlation of treated areas are given in Supplement A.
The posterior distribution of the model parameters and the missing control potential outcomes for the treated units is derived using Hamiltonian Monte Carlo (HMC) implemented in Stan (Carpenter et al., 2017) and its R interface (Stan Development Team, 2018). The code developed to obtain the results presented in this paper is publicly available at the following GitHub repository: https://github.com/Giulio-Grossi/SVR. This software enables the replication of all experiments described in the paper and it can be employed in any setting where a generic distance measure between units is defined. As a result, the proposed method is highly flexible and can be adapted to a variety of applications and analytical contexts. Furthermore, the repository includes the code for obtaining estimates and confidence intervals using the other methods evaluated in the simulations of Section 6.
@@ 6.1 Data generating process
In this section, we present a simulation study designed to assess how spatial correlation among units affects the regression coefficients and to evaluate the performance of our method relative to the original SC approach and a standard vertical regression under various scenarios, including a benchmark scenario with no spatial structure.
We consider a causal panel study with $N_{1}=5$ treated areas and $N_{0}=10$ control areas observed in $T_{0}$ pre-treatment periods and $T-T_{0}$ post-treatment periods. We consider three lengths of the pre-treatment period, $T_{0}=10,20,40$, combined with the following lengths of the post-treatment period: $T-T_{0}=5,10,T_{0}/2$, for a total of seven distinct scenarios (for $_{0}=10, TT$$T_{0}=T_{0}/2=5$, and for $_{0}=20, TT$$T_{0}=T_{0}/2=10$). Therefore, we consider simulation scenarios closely aligned with the empirical setting described in Section 3. In particular, we consider a short panel ($T_{0}=10$) with a similar number of treated and control units as in our study, and allow for different levels of spatial correlation to reflect empirically plausible spatial dependence structures.
@@
We set $\sigma^{2}_{s}=0.4$ and consider four values of the lengthscale parameter $\rho_{s}^{2}$: $0.001^{2}$, $0.2^{2}$, $0.4^{2}$, and $0.6^{2}$ corresponding to scenarios where the weights associated with the same control unit for different treated units are essentially independent, weakly correlated, mid-correlated, and strongly correlated, respectively. These scenarios mimic studies where accounting for the spatial correlation structure among the treated units in estimating the regression weights may be progressively more relevant, and the case of $\rho_{s}^{2}=0.001^{2}$ represents a scenario where our approach is not expected to provide any advantages. In Supplement B, we include a visualization of $\bm{\beta}_{c}$ in our simulations under the different values of $\rho_{s}^{2}$.
We generate temporally-correlated outcomes for the control areas. Specifically, let $\bm{Y}_{c}^{\bm{z}_{0}}=\bm{Y}_{c}=(Y_{c1},\ldots,Y_{cT})^{\top}$ represent the times-series under no intervention for the $c^{th}$ unit for $c=N_{1}+1,\ldots,N_{1}+N_{0}$, which we observe. We draw $\bm{Y}_{c}$ from a $T$-dimensional multivariate normal distribution. We specify mean vector $t_{c}\bm{1}_{T}$ where $t_{c}\sim\mathcal{N}(0,0.7^{2})$. The covariance matrix is set to $\Sigma_{T\times T}=\sigma^{2}_{c}\mathcal{K}(D_{T},\rho^{2}_{c})+\eta_{c}\mathbf{I}_{T}$, where $D_{T}$ is a $T\times T$ matrix of time differences and $\mathbf{I}_{T}$ is the $T\times T$ identity matrix, so that
$\operatorname{Cov}(Y_{ct},Y_{ct^{\prime}})=\sigma^{2}_{c}\operatorname{exp}\left(-\frac{[D_{T}]_{tt^{\prime}}}{2\rho_{c}^{2}}\right)+\eta_{c}\mathrm{I}\{t=t^{\prime}\},$
and $\mathrm{I}\{\cdot\}$ is the indicator function. We set $\sigma^{2}_{c}=0.3^{2}$, $\rho^{2}_{c}=0.05^{2}$, $\eta_{c}=0.15^{2}$, for all $c=N_{1}+1,\ldots,N_{1}+N_{0}$. In Supplement B, we illustrate that the outcome time series for control units under this design has temporal correlation that resembles the data in Figure 2.